In how many ways could a committee of five people be selected from eight females and six males if the committee must have more females than males and each gender must be represented.
Answer by jim_thompson5910(35256)

Note: I'm using the formula n C r = (n!)/(r!(n-r)!)


Break it down into cases

4 females, 1 male: (8 C 4)*(6 C 1) = 70*6 = 420

3 females, 2 male: (8 C 3)*(6 C 2) = 56*15 = 840

420+840 = 1,260


There are 1,260 ways to form the committee.